OpenStudy (anonymous):

I need help with Logarithms... I freaking hate them! ;-; someone help please?

OpenStudy (anonymous):

uhm... tried that place.

OpenStudy (anonymous):

it was the first place I looked.

OpenStudy (anonymous):

but thanks anyways @Pandabear500

OpenStudy (anonymous):

anyone else willing to help?

hartnn (hartnn):

do you have a specific question you need help with?

OpenStudy (anonymous):

no more like I wanna understand them in general

OpenStudy (anonymous):

I am freaking crying over these cause of my school work and I want to learn ways to be able to solve them without going through a ton of freaking work

OpenStudy (anonymous):

@hartnn you there dude? or are you looking up something?

hartnn (hartnn):

There are just a bunch of formulas you need to practice while studying Logarithm.. Also, if Khan Academy doesn't help, I am not sure what will! maybe mathisfun.com ?

OpenStudy (anonymous):

oh. thanks anyways.

OpenStudy (mrnood):

It's difficult to help when you don't have a specific issue. "Logarithms" is a large topic and I can't write the book on them here. Have you studied the rules of 'exponents' prior to this - because logs are based on those rules (e.g. a^m * a^n = a ^ m+n) etc.??

OpenStudy (anonymous):

I know the formulas. its executing them that I have a problem with

OpenStudy (mrnood):

mostly using logs is just addition or subtraction , sometimes multiplication and division. But the whole point of logs is to use addition of exponents etc. instead of multiplication... Do you understand that adding logs is adding exponents which is ht same as multiplying?

OpenStudy (anonymous):

or subtracting right?

OpenStudy (mrnood):

subtracting logs is subtracting exponents which is the same as dividing...

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

but could you give me an example so I can take notes?

OpenStudy (anonymous):

I think his computer either froze or he got disconected

OpenStudy (mrnood):

I can - but htere are plenty of better texts and sites available with examples. Nobody can implant the knowledge in oyu without 'freakin' work. Here is a very simple example: 100 * 1000 = x 10^2 * 10^3 = x take logs: log 10^2 + log 10^3 = log x 2 + 3 = log x 5 = log x 10^5 = x 100,000 = x

OpenStudy (anonymous):

ah alright and I deserved that ^^;

OpenStudy (anonymous):

thanks! that actually helped a lot. I have no idea what I was so worried with.

OpenStudy (mrnood):

np